# Semi-Centennial REDUCE

### v1

@@
Abstract.
We present a version of the reduce computer algebra system as it was in the early 1970s. We show how this historical version of reduce may be built and run in very modest present-day environments and outline some of its capabilities.
copyright: nonejournalyear: 2025conference: ISSAC Software Demonstrations; July 28–Aug 1; Guanajuato, Mexico
1. Introduction
@@
We note that even the most meager of today’s common computing platforms are significantly more powerful than the hardware on which this version of reduce was deployed. One may therefore contemplate how significant computer algebra code can be incorporated in small devices such as micro-controllers. This takes us further into the territory explored by the Compact Computer Algebra workshop series, e.g. (Fitch09, ).
@@
Both authors have a historical connection with reduce, though to greatly different degrees. One of us (ACN) has been one of the principal contributors to the system over almost its entire existence, e.g. (Norman93, ). The other of us (SMW) was a user of reduce in the 1970s for general relativity computations: after performing the first of two similar lengthy computations by hand over the course of three months (on the advice of a supervisor), it was possible to learn reduce and complete the second computation in one weekend (watt81, ).
@@
In what follows, we summarize some of the key steps that led to this early version of reduce, outline the issues encountered in building a working version from archival sources, and outline some of its capabilities. This is a companion to the ISSAC paper (Corless+:25, ) where a live demonstration of this historical reduce system will be made.
2. Early REDUCE
@@
In 2005 Tony Hearn, the originator of the reduce algebra system, presented a paper “reduce - The First Forty Years” (Hearn:05, ) explaining that in the 1960s he was working on calculations involving Feynman Diagrams and that John McCarthy suggested he have a go at using the new language Lisp to mechanize some of the really tedious work, and also offered him access to what was at the time a state of the art large computer suitable for such a project. By the start of 1970 the code that had originated as specifically aimed at High Energy Physics calculation had extended to become suitable for a range of uses. And in 1971 he presented this to a fully international audience (Hearn:71a, ). Some important papers leading up to this were (hearn66, ; hearn68, ; hearn69, ). Building on collaboration with people he met at that and other conferences he prepared a test and demonstration script for reduce that included samples of calculations relevant to celestial mechanics and general relativity as well as high energy physics and rather general basic computation. In 1985 Marti and Hearn collected timings for a wide range of machined and reported on “REDUCE as a Lisp Benchmark” (Marti:85, ). A version of that test script is part of the set of regression tests shipped with the system in its open source form even today.
@@
Fairly recently some archive tapes from the early 1970s were recovered and decoded, and found to contain a set of reduce sources from 1972-73. Inspection shows that the user interface they offer has hardly changed in the intervening 50 years and that quite large segments of the code are utterly recognizable save for a conversion from all upper case to lower case lettering. The version of the test script included there differs in a very few rather minor ways from the one used today. It was possible to build and run this historic version of the software.
@@
The version we have available for demonstration now sits on a naively-implemented and fairly minimal Lisp system coded in C that uses just an interpreter (i.e. no compiler) and so by today’s standards the performance there is severely compromised. However on even cheap machines today such as the Raspberry Pi we can compare its absolute timings with a wide range of systems from the past that were reported in the benchmark paper previously cited. The times we see are well up with mainframes from earlier times and with specialized workstations such as dedicated Lisp machines and the amount of store we can now deploy is absurdly greater than any of the early-days options.
@@
By reviewing this version of the system from the early 1970s we can see the growth in capability (and bulk of source code) that development with multiple international contributors has grown reduce to, but also be reminded how many of the essential core capabilities that if offers were available that far back. And we can be reminded of the impact of Moore’s Law and realize just what a challenge getting anything going well will have been that long ago.
3. Building The System Today
The files used mostly derive from a version of the reduce source from 1973 as recovered from some archive files. The comments here are by ACN reporting on the experience of getting the code to run.
@@ 3. Building The System Today
The version started with is the file "reduce.red" which was for use on TENEX (on a PDP10) and is dated 25 March 1973. The process of making this run will have been very similar to the porting process normally needed for reduce then when Lisp systems on the various machines it was used on differed. The start of the file contains a collection of mappings, definitions and redefinitions that Lisp 1.6 on TENEX did not support or where it used different names. It was necessary to replace or override these with code that has the same intent but to set up compatibility with the small Lisp system used.
The code shows signs of a range of compatibility issues it had encountered. The ones that seem most notable are enumerated below. There is also a reduce source file dated 15 September 1972 that is all in Lisp. It includes many checks of the form (EQ *MODE (QUOTE SYMBOLIC) and that amounts to a signature of a system where executable parts of reduce would be written in its own language. As in item (1) that follows it is clear that LISP/360 was not offering a string data type. A version of reduce in its own language needed a directly Lisp-coded for bootstrapping, so Lisp versions like this will have persisted for a while even when the main development had moved on.
  1. (1)
@@
  2. (2)
    There are functions EXPLODE and COMPRESS that map between symbols and numbers (and perhaps strings) and lists of characters. The behaviour of these on symbols whose names included whitespace or punctuation marks was not fully stable. Specifically using COMPRESS on a list consiting of (* A B C) might produce a symbol called *ABC or it might stop at the “*” and need escape characters to support unusually spelt characters. And then EXPLODE might or might not insert those escape characters.
  3. (3)
@@
       (X := LSH (MAKNUM(CAAR GET(X,’PNAME),’FIXNUM),-11))>64
@@
    where the MAKNUM and the shift are not very portable and it is clear that lower case letters had not been invented.
  8. (8)
@@
  9. (9)
    Anything that is specific to iterative vs batch use is of course “delicate”.
The PDP10 scheme for covering all the above is just a clearly separated out 500 line prelude to the main reduce code.
@@
Once reduce is built it should be able to parse its own source - but of course at the start of testing it has not been built! So to bootstrap it one starts by using an existing running version of reduce to parse the file "reduce.red" and emit the raw Lisp equivalent. Well of course for pragmentic reasons some of the machine-sensitive customizations are done ahead of that conversion. A current version of reduce was used to make a file "reduce.lsp". For this conversion a number of behaviours that the current parser would have exhibited needed to be disabled, either by removing tags from property lists or by hiding some symbols completely. The file "reduce.red" was then hand-edited further to gain the level of compatibility between 1973 and 2025 standards that was needed.
@@
The resulting reduce.lsp was loaded into a Lisp system known as ”vsl”. This is a fairly minimal Lisp that consists of 3500 lines of C code implementing much of the ”Standard Lisp” that later versions of reduce settled on. It is complete enough that it can run an up to date version of reduce, but is an interpreter-only system and so is not very fast. The aim was that the C code should be fairly spartan, and so for instance at that level it supports CAR and CDR. All the combinations such as CDADR and so on are then defined in Lisp in a file "vsl.lsp". That file also provides the functions to append and reverse and search lists. But the two big sections of codin in vsl.lsp are the ones that support big integers and that can prettyprint (ie print with neat indentation) Lisp code. Well in 1973 many Lisp systems did not support integers with more bits than would fit in a machine register. vsl.c provides for a represention of overlarge integers but leaves all arithmetic on them to lisp code in vsl.lsp, where the numbers are handled as list of digits. Performance took second place to code simplicity. It would have been possible to modify vsl to support the ancient reduce more directly, but instead "reduce.lsp" was modified to live with it.
@@
The result of all of the above is a workable reduce (which will without doubt have residual errors and limitations above and beyond the ones inherent from its age). The test applied for it is a file that in a modern copy of reduce would live as packages/alg/alg.tst and which was for very many years treated as the definitive demonstration and main test for reduce. Tony Hearn collected timings from running it on many computers (dedicated Lisp machines, departmental servers, mainframes, supercomputers and various more personal machines) and today it feels amazing that the really crude implementation reported here is dramatically faster then a whole range of impressive and expensive hardware from the past (Marti:85, ). As what may be a key illustration of this test ran for most of a minute on a VAX 11/780 or for 5 or 6 seconds on the fastest IBM mainframe of the day. The current crudely implemented interpreter-only Lisp running on a 7 year old (and hence not really cutting edge) x86_64 pc completes the run in about 0.3 second.
@@
Note that the underlying representation used for polynomials and for rational functions has conspicuously not changed from usage in the early sources. The functions ADDF, ADDSQ and flags such as !*MCD, along with the general framework for simplifications using a function SIMP that dispatches based on the identity of the expression it is handed has the same overall structure even though a range of detailed changes and upgrades have been applied. The same holds true for much of code for both parsing and printing: the function names in the old code will be really familiar to those who work with the current system.
4. A Sample of the System Capabilities
@@
This version of reduce was fully capable of polynomial and rational function arithmetic and computations involving elementary and unknown functions.
To be compatible with the 1973 version of reduce, it was necessary to modify the current standard test file "alg.tst" in some minor syntactic ways:
  • to replace the use of <<...>> blocks with BEGIN ... END,
  • to unwind the use of WHERE into LET, and
  • to replace the result variable "ws" with "!*ans".
With these modifications, the test file shows:
@@
  • solution of the $f$ and $g$ series problem (Sconzo65, ),
@@
  • solution to a problem from the textbook Relativistic Quantum Mechanics (Bjorken+64, ).
The test file for the 1973 reduce is shown in the appendix.
5. Conclusions

### v3

@@
Abstract
We present a version of the reduce computer algebra system as it was in the early 1970s. We show how this historical version of reduce may be built and run in very modest present-day environments and outline some of its capabilities. This shows that a large part of computer algebra function can be delivered by a system that is small by today’s standards and can be deployed in resource-constrained settings.
I Introduction
@@
We note that even the most meagre of today’s common computing platforms are significantly more powerful than the hardware on which this version of reduce was deployed. One may therefore contemplate how significant computer algebra code can be incorporated in small devices such as micro-controllers. This takes us further into the territory explored by the Compact Computer Algebra workshop series, e.g. [4].
@@
Both authors have a historical connection with reduce, though to greatly different degrees. One of us (ACN) has been one of the principal contributors to the system over almost its entire existence, e.g. [11]. The other of us (SMW) was a user of reduce in the 1970s for general relativity computations: after performing the first of two similar lengthy computations by hand over the course of three months (as required by a thesis advisor), it was possible to learn reduce and complete the second computation in one weekend [14].
@@
In what follows, we summarize some of the key steps that led to this early version of reduce, outline the issues encountered in building a working version from archival sources, summarize some of its capabilities and discuss its performance. This 1973 version of the software is publicly available for download [3]. The software portability issues encountered in its development are recounted in [12] and the historical context is given in [2].
II Early REDUCE
@@
In 2005 Tony Hearn, the originator of the reduce algebra system, presented a paper “reduce — The First Forty Years” [9], explaining that in the 1960s he was working on calculations involving Feynman Diagrams and that John McCarthy suggested he have a go at using the new language Lisp to mechanize some of the really tedious work, and also offered him access to an IBM 7090. At the time, this was a state of the art large computer suitable for such a project. By the start of the 1970s, the code that had originated as specifically aimed at High Energy Physics calculations had extended to become suitable for a range of uses. In 1971 he presented this work to a fully international audience [8]. Some important papers leading up to this were [5, 6, 7]. Building on collaboration with people he met at that and other conferences, he prepared a test and demonstration script for reduce that included samples of calculations relevant to celestial mechanics and general relativity as well as high energy physics and rather general basic computation. In 1985 Marti and Hearn collected timings for a wide range of machines and reported these in “reduce as a Lisp Benchmark” [10]. A version of that test script is part of the set of regression tests shipped with the system in its open source form even today.
@@
Fairly recently some archive tapes from the early 1970s were recovered and decoded, and found to contain a set of reduce sources from 1972–73. Inspection shows that the user interface they offer has hardly changed in the intervening 50 years and that quite large segments of the code are completely recognizable, with only a conversion from all upper case to lower case lettering. The version of the test script included there differs in a very few rather minor ways from the one used today. It was possible to build and run this historic version of the software.
@@
The version we have recovered and made available for demonstration now sits on a naively-implemented and fairly minimal Lisp system coded in C that uses just an interpreter (i.e. no compiler) and so by today’s standards the performance there is severely compromised. However, on even cheap machines today such as the Raspberry Pi, we can compare its absolute timings with a wide range of systems from the past that were reported in the benchmark paper previously cited. The times we see are similar to those on mainframes from earlier times and with specialized workstations such as dedicated Lisp machines and the amount of store we can now deploy is absurdly greater than any of the early-days options.
@@
By reviewing this version of the system from the early 1970s we can see the growth in capability (and bulk of source code) that development with multiple international contributors has grown reduce to, but also be reminded how many of the essential core capabilities that it offers were available that far back. And we can be reminded of the impact of Moore’s Law and realize just what a challenge getting anything going well would have been that long ago.
III Building The 1973 System Today
The files we have used are available for download [3]. They derive mostly from a version of the reduce source from 1973 as recovered from some archive files. This version was for TENEX on a PDP 10 but we also have sources dated September 1972 where the comments make it clear that it was for use on the MTS system running on an IBM 360. The PDP10 version was slightly more complete and easier to work from.
@@ III Building The 1973 System Today
The main reduce source code is found in the file "reduce.red". The process of making this run will have been very similar to the porting process normally needed for reduce then when Lisp systems on the various machines it was used on differed. The start of the file contains a collection of mappings, definitions and redefinitions that Lisp 1.6 on TENEX did not support or where it used different names. It was necessary to replace or override these with code that has the same intent but made compatible with the small Lisp system used.
The code shows signs of a range of compatibility issues it had encountered. The ones that seem most notable are enumerated below. The PDP10 scheme for covering all of this is just a clearly separated out 500 line prelude to the main reduce code.
  1. 1.
@@
  2. 2.
    There are functions EXPLODE and COMPRESS that map between symbols and numbers (and perhaps strings) and lists of characters. The behaviour of these on symbols whose names included whitespace or punctuation marks was not fully stable. Specifically using COMPRESS on a list consisting of (* A B C) might produce a symbol called *ABC or it might stop at the “*” and need escape characters to support unusually spelt characters. And then EXPLODE might or might not insert those escape characters.
  3. 3.
@@
       (X := LSH (MAKNUM(CAAR GET(X,’PNAME),
                         ’FIXNUM),-11))>64
@@
    where the MAKNUM and the shift are not very portable and it is clear that lower case letters were not been considered.
  8. 8.
@@
  9. 9.
    Anything that is specific to interactive vs batch use is of course “delicate”.
Once reduce is built it must be able to parse its own source — but of course at the start of testing it has not been built! So to bootstrap it one starts by using an existing running version of reduce to parse the file "reduce.red" and emit the raw Lisp equivalent. In the files, there is also a reduce source file dated 15 September 1972 that is all in Lisp. It includes many checks of the form (EQ *MODE (QUOTE SYMBOLIC)) and that amounts to a signature of a system where executable parts of reduce would be written in its own language. A version of reduce in its own language needed a part directly Lisp-coded for bootstrapping, so Lisp versions like this will have persisted for a while even when the main development had moved on.
@@
For pragmatic reasons some of the machine-sensitive customizations are done ahead of that conversion. A current (i.e. 2025) version of reduce was used to make a file "reduce.lsp". For this conversion a number of behaviours that the current parser would have exhibited needed to be disabled, either by removing tags from property lists or by hiding some symbols completely. The file "reduce.red" was then hand-edited further to gain the level of compatibility between 1973 and 2025 standards that was needed.
@@
The resulting "reduce.lsp" was loaded into a Lisp system known as ”vsl”. This is a fairly minimal Lisp that consists of 3500 lines of C code implementing much of the ”Standard Lisp” that later versions of reduce settled on. It is complete enough that it can run an up to date version of reduce, but is an interpreter-only system with no optimizations and so is not at all fast. The aim was that the C code should be fairly spartan, and so for instance at that level it supports only CAR and CDR. All the combinations such as CDADR and so on are then defined in Lisp in the file "vsl.lsp". That file also provides the functions to append and reverse and search lists. But the two big sections of code in "vsl.lsp" are the ones that support big integers and that can prettyprint (i.e. print with neat indentation) Lisp code. In 1973 many Lisp systems did not support integers with more bits than would fit in a machine register. The file "vsl.c" provides for a representation of overlarge integers but leaves all arithmetic on them to lisp code in "vsl.lsp", where the numbers are handled as list of digits. Performance took second place to code simplicity. It would have been possible to modify vsl to support the ancient reduce more directly, but instead "reduce.lsp" was modified to live with it. We note that "vsl.c" uses zlib compression so that saved heap images consume less space and histedit to support more civilised terminal interface. Neither of these is intrinsically necessary for reduce – they are just used in the small Lisp system to make life easier.
@@
The result of all of the above is a workable reduce (which will without doubt have residual errors and limitations above and beyond the ones inherent from its age). The test applied for it is a file that in a modern copy of reduce would live as packages/alg/alg.tst and which was for very many years treated as the definitive demonstration and main test for reduce. Tony Hearn collected timings from running it on many computers (dedicated Lisp machines, departmental servers, mainframes, supercomputers and various more personal machines) and today it feels amazing that the really crude implementation reported here is dramatically faster then a whole range of impressive and expensive hardware from the past [10].
@@
Note that the underlying representation used for polynomials and for rational functions has importantly not changed from usage in the early sources. The functions ADDF, ADDSQ and flags such as !*MCD, along with the general framework for simplifications using a function SIMP that dispatches based on the identity of the expression it is handed has the same overall structure even though a range of detailed changes and upgrades have been applied. The same holds true for much of the code for both parsing and printing: the function names in the old code will be really familiar to those who work with the current system.
IV The System Capabilities in Brief
@@
This 1973 version of reduce was fully capable of polynomial and rational function arithmetic and computations involving elementary and unknown functions, including user-defined simplification rules.
Reverting a few minor syntax changes, described below, the current standard reduce test file, "alg.tst", is able to run in the 1973 system. That file shows:
@@
  • solution of the $f$ and $g$ series problem [13],
@@
  • solution to a problem from the textbook Relativistic Quantum Mechanics [1].
The contents of the test file for the 1973 reduce is given in the appendix. To be compatible with the 1973 version of reduce, it was necessary to modify the current standard test file "alg.tst" in [3] to older syntax as follows:
  • replace <<...>> blocks with BEGIN ... END,
  • unwind the use of WHERE into LET, and
  • replace the result variable ws with !*ans.
Despite its early appearance and small size, reviewing various test and demonstration files that came with reduce and the reports on applications in published work citing it, suggests that it could address a broad swathe of problems. Users would often need to re-cast their problems to allow for its restrictions (e.g. notably avoiding need for the calculation of polynomial GCDs wherever possible, for instance by introducing new symbols to stand for more complicated expressions that might appear in denominators). And where there were intrinsic limitations (such as a lack of floating point arithmetic or support for big integers) those were in fact addressed over the next few years.
V Observations
We can compare the 1973 version of reduce and its computing environment to those of today. The source is a little over 6000 lines long (170 Kbytes) as against the current version which is over 500,000 lines and almost 19 Mbytes. Note that Reduce 3.8, a version from 2007 fairly shortly before the code was released as open source, was 300K lines and 10 Mbytes.
[…]
IBM Raspberry
370-168 model 1 Pi Zero 2W
Year Introduced 1972 2021
Cost (usd 2024) 35–50,000,000 15
Size (m2) 90-140 0.002
Power w/o periph. (W) 50–70,000 0.6
Cooling active water passive air
Instruction Set CISC RISC
Clock Speed (MHz) 2 2400
Word Size (bits) 32 64
Primary Mem. (MB) 1–8 512
Secondary Mem. (MB) 200 64,000 (typical)
[…]
TABLE I: High-end 1973 computer vs low-end 2025 computer
With this in mind, we can compare the time to run the "alg.tst" test file on various historical computer models versus the time to run it with the 1973 version of reduce on modern hardware. A few timings are shown in Table II, where some notable historical times are taken from [10]. The DEC KL-10 was the machine on which macsyma was made available over the ARPANET, the VAX 11/780 was the machine on which most of Maple’s initial development took place, the IBM 308X series was the platform on which the scratchpad ii (later named axiom) had most of its initial development, and the Symbolics 3600 was a single-user computer optimized for Lisp and featuring macsyma. The Xerox Dolphin is representative of the early years and the largest memory of these historical machines reported was the Symbolics 3600, which could be configured with up to 7.5 megawords of memory (around 32 Mbytes). The VAX 11/780 and IBM 3084 were the mainstays of academic use in their time.
We demonstrated the 1973 version of reduce running on a Raspberry Pi Zero at ISSAC 2025. Rather than using a modern, optimised, full-scale Lisp system in this demonstration we chose to build Reduce on top the naïve “vsl” Lisp system as described earlier, and whose C source is included with the historical reduce.
When we compare reduce speed on the cut-down Lisp with a version running on a production-quality Lisp, we see a speed factor of about 6, so current computers, even tiny ones, are even faster as compared with the biggest ones of the past than the figures we have presented might initially suggest.
VI Conclusions
@@
Model CPU time (s)
Contemporary reduce versions on then state-of-the-art Lisps
Xerox Dolphin 322.0
DEC KL-10 24.6
DEC VAX 11/780 50.3
Symbolics 3600 45.0
IBM 370 158 49.9
IBM 3084 5.4
1973 reduce on the naïve Lisp described in Sec III
Raspberry Pi Zero 2W 4.416
Raspberry Pi 5 0.599
MacBook Pro (M3 Max) 0.239
Modern reduce on Cambridge Standard Lisp
Raspberry Pi 5 0.099
MacBook Pro (M3 Max) 0.050
TABLE II: Time to run "alg.tst". Lower is better.